OpenStudy (anonymous):

Help!! Evan is painting a stool that has a cylindrical seat and four rectangular prism legs. The seat has a diameter of 14 inches and a height of 4 inches. Each leg is 4 inches by 4 inches by 18 inches. How much area will Evan have to paint? Use pi for 3.14 Show your work.

OpenStudy (anonymous):

sa of a cylinder formula is 2 π r2 + 2 π r h sa of a rectangular prism formula is Find the area of two sides (Length*Height)*2 sides Find the area of adjacent sides (Width*Height)*2 sides Find the area of ends (Length*Width)*2 ends \[2(3.14)7^{2}=307.72\] \[2(3.14)(7)(4)=175.84\] 175.84+307.72=483.56 4x4=16 4x18=72 4x18=72 72+72+16=160 4(160)+483.56=1123.56

OpenStudy (anonymous):

inches squared

OpenStudy (anonymous):

Where didi you get the 4 from

OpenStudy (anonymous):

4 legs

OpenStudy (anonymous):

Oh okay

OpenStudy (anonymous):

Also could you help me on some other ones if its okay

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

A rectangular prism has a volume of 700 cubic centimeters, and the base has an area of 25 square centimeters. What is the height of the prism? 28 36 45 56 The lateral area of a cube is 576 square kilometers. What is its side length? 6 12 24 36

OpenStudy (anonymous):

the first answer is 28 you do 700/25

OpenStudy (anonymous):

\[LA=4s ^{2}\] the second is 12 you do 576/4 which is 144 then take the square root which is 12

OpenStudy (anonymous):

Okay thanks so much

OpenStudy (anonymous):

You're Welcome